A hundred grams of sulfuric acid yielded ten grams of water.



1
Expert's answer
2021-10-14T04:05:49-0400

Balanced

H2SO4(l) = H2O(l) + SO3(l)


Moles of H2SO4 = "\frac{100}{98}=1.02 moles"


Moles of H2O "=1.02 moles"


Not applicable for limiting and excess reactant


Theoretical yield = 1.02 × 18=18.36 g


Percent yield = "\frac{18}{18.36}×100=98" %